OpenStudy (anonymous):

GRAPH THE LINEAR SYSTEM AND ESTIMATE THE SOLUTION. THEN CHECK THE SOLUTION ALGEBRAICALLY.

OpenStudy (anonymous):

3X+2Y=-4 X+3Y=1

OpenStudy (anonymous):

please explain.

OpenStudy (ikram002p):



OpenStudy (ikram002p):



OpenStudy (ikram002p):

my sketch is not accurate but the idea is the point of intersection between the two lines is the solution

OpenStudy (ikram002p):

ALGEBRAICALLY. use the too equation to find x , y \[7x=-13\rightarrow x=\frac{ -7 }{ 13},y=\frac{ 98 }{ ? }\] So the solution would be \[\left( \frac{ -7 }{ 14} ,\frac{ 98 }{ 13 }\right)\] the point that intercect must be close to this point .

OpenStudy (anonymous):

im confused on how you graphed them.

OpenStudy (anonymous):

did you turn in into y=mx+b form